OpenStudy (anonymous):

How many iron atoms are there in 2.50moles of each of the following? FeWO4 = ? atoms of Fe Could someone explain this step by step???

sam (.sam.):

FeWO4?

OpenStudy (anonymous):

yes, Fe W O4

OpenStudy (anonymous):

wat is W>>>?

OpenStudy (anonymous):

w is tungsten with atomic weight of 183.84

OpenStudy (anonymous):

Ok...... No : of moles = Mass / Molar mass

OpenStudy (anonymous):

1 st find the mass..

OpenStudy (anonymous):

The mass of FeWO4 I got was 303.69g, and for the Fe I have 55.85

OpenStudy (anonymous):

FeWO4 = 56 + 183+ 64 = ? = Molar Mass

OpenStudy (anonymous):

Ok....Find mass

OpenStudy (anonymous):

Isnt that the mass..?

OpenStudy (anonymous):

That is molar Mass

sam (.sam.):

So, Iron (II) tungstate \[2.5moles~FeWO_4 \times \frac{6.02 \times 10 ^{23} atoms~of~FeWO_4}{1mole~of~FeWO_4} =1.51 \times 10^{24} \\ \\ 1.51 \times 10^{24} \text{ Multiplied by 7 since there's 7 atoms}\\ \\ =1.05 \times 10^{25} atoms\]

OpenStudy (anonymous):

that looks beautiful. but.. where did you get the seven from?

OpenStudy (anonymous):

2.50 * 303.69= 759.225 = mass

sam (.sam.):

There are 7 atoms in FeWO4

OpenStudy (anonymous):

@.Sam. the question asks for no of Fe atom

OpenStudy (anonymous):

@.Sam. I think there's 6?? 1 +1+4?? and @Yahoo! Okay, I got to there as well... how what?

OpenStudy (anonymous):

now* oh and thank you guys for your help!

OpenStudy (anonymous):

Do u have answer key with u

OpenStudy (anonymous):

No, I don't..

OpenStudy (anonymous):

i am sure,,,that sam.....got it..wrong..)

OpenStudy (anonymous):

My answer is 8.18 * 10^24 atoms

OpenStudy (anonymous):

How did you get there?

OpenStudy (anonymous):

is the first step 2.50moles of FeWO4 x 6.022 x 10^23 particles of FeWO4/ 1mole

OpenStudy (anonymous):

Silvermouse has it - you divide by 1 mole of FeWO4 so you can multiply by Avagadro's number (6.022 * 10^23) that will get you units of *atoms, or *particles, *formula units it a set conversion number to determine any of those units I've *listed.